OpenStudy (anonymous):

Question 9. 9. Select the pair of equations whose graphs are perpendicular. (Points : 5) 5x – 8y = 9 12x – 5y = 7 x + 6y = 8 y = 2x – 8 y = 2x – 7 x + 2y = 3 2y = –3x + 5 2x + 3y = 4 Question 10. 10. Which is the equation of the line passing through (0, 6) and (3, 0)? (Points : 5)

OpenStudy (anonymous):

When you multiply the slopes of the two lines you should get -1. So the third equation appears to be the correct answer. y= 2x-7 So slope is 2 x+ 2y = 3 or y= -x/2 +3 So slope is -1/2 (-1/2)(2) = -1